OpenStudy (anonymous):

Last 1 for the night i PROMISE JUST CHECK :)

OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, can be written as 7sqrt(a).

OpenStudy (anonymous):

My answer= sqrt(144x^4) Simplified 12x^2

OpenStudy (anonymous):

@Callisto @Smoothmath :D

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

i sound like i'm 6 lol

OpenStudy (anonymous):

The greatest perfect squares method is mostly used for things that AREN'T already perfect squares. You picked 144 because you can just square root it. That's not really the greatest perfect squares method. The idea with the greatest perfect squares method is that the number under the root ISN'T a perfect square, and so we go through and find out if it has any factors that ARE perfect squares.

OpenStudy (anonymous):

A super simple example is sqrt(8) To simplify that, the first thing that I notice is that it ISN'T a perfect square, so I can't just square root it. So I start to check if there are any perfect squares that are factors of 8. 4 is a perfect square. Does it go into 8? Yes. Sqrt(8) = sqrt(4*2) = sqrt(4)*sqrt(2) = 2*sqrt(2)